OpenStudy (anonymous):

Simplifly the expression : √39 (√6 + 7)

ganeshie8 (ganeshie8):

before multiplying √39 thru out the expression, write 39 as 13x3 and 6 as 3x2

ganeshie8 (ganeshie8):

\(\large \sqrt{39}(\sqrt{6} + 7)\) \(\large \sqrt{13 \times 3}(\sqrt{3 \times 2} + 7)\) \(\large \sqrt{3^2 \times 26 \times} + 7\sqrt{39}\)